OpenStudy (anonymous):

5x−8y=15 What is the x-intercept?

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

5x−8y=15 5x−8*0=15 5x + 0 = 15 5x = 15 x = 15/5 x = 3 So the x-intercept is the point (3,0). This is where the graph crosses the x axis